A traffic study of 2,500 vehicles that passed by a checkpoint showed that their speeds were normally distributed with a mean of 82.5 kph and a standard deviation of 11.5 kph. How many vehicles had a speed of more than 90 kph?



1
Expert's answer
2021-11-12T13:35:07-0500

"\mu=82.5 \\

\sigma= 11.5 \\

P(X>90) = 1 -P(X<90) \\

= 1 -P(Z< \frac{90-82.5}{11.5}) \\

= 1 -P(Z< 0.652) \\

= 1 -0.7428 \\

= 0.2572 \\

N = 2500 \times 0.2572 = 643"

Answer: 643